Other formats

    Adobe Portable Document Format file (facsimile images)   TEI XML file   ePub eBook file  


    mail icontwitter iconBlogspot iconrss icon

Victoria University Antarctic Research Expedition Science and Logistics Reports 1980-81: VUWAE 25

Event 14 - McMurdo Sound bottom sampling programme

Event 14 - McMurdo Sound bottom sampling programme.

Oct. 9 Ward, Pyne and Fitzgerald to Christchurch.
12 Ward, Pyne, Fitzgerald and Garrick to Scott Base after two days delay in Christchurch.
13-14 Preparing and checking field gear, scientific equipment and vehicles (Toboggans).
15 A full day spent by the whole party drilling and blasting the ice quarry for Scott Base water supply.
16-18 Continued with field equipment preparation, testing toboggans and sledges on an overnight survival course (17-18).
18 Ward, Pyne and Scott Base Leader (Clark) on helo reconnaissance of the sea-ice conditions during an afternoon training flight to Cape Royds, Butter Point and Cape Chocolate.
19-21 Remainder of the field equipment checked. Ice recon. photographs developed.
22 Testing the coring device at McMurdo Station Desalination effluent site. Corer malfunctioned, returned to Scott Base in the evening.
23 Repairs to both corer (butterfly valve) and toboggan executed.
24 Ward, Pyne, Fitzgerald and Garrick departed Scott Base for Cape Evans (morning). Camp established on the ice edge at Evans (North Bay). Evening reconnaissance to the ice edge at Cape Royds, returning to camp three hours later.
25-28 Tent days. Southerly storm system; high winds and blizzard conditions with blowing snow and often zero visibility.
28 "Broke camp" and returned to Scott Base; arrived early morning.
(evening) Difficulties were only encountered between the Dellbridge Is. on
-29 bare ice in high winds.
29-Nov. 1 Repairs to toboggans and refitting of new sledges at Scott Base.
2 Sampling in front of McMurdo Station at the Desalination Effluent outlet.
3 Modifying corer at Scott Base to add extra 60 lbs weight.
4 Coring attempt at McMurdo resulted in a small core (in hard ground). Moved to old Fish Hut site; 5km from McMurdo, 476m water depth. Unsuccessful attempt, with repaired valve again broken.
5-6 New valve, flown from N.Z. fitted to corer and final preparations made to sledges and toboggans.
7 Departed Scott Base at 9.00 hours for Butter Point on a direct route to tip of "Dirty Ice" then to Butter Point. A strip of rough ice negotiated 2-6km off the Bowers extending from Strand Moraines to Butter Point. Arrived Butter Point 1800 hours, established camp at the MSSTS refuge hut.page 61
Nov. 8 Moderate winds prevented work in the morning. Two holes attempted in late afternoon; corer malfunctioned on hole 2 about 5km east of Butter Point.
9 Corer repaired and another coring attempt later in the day. Trigger mechanism failure on the corer.
10 Corer repaired during the day. Neale (surveyor) arrived by helo from Vanda early evening. Ward, Pyne, Fitzgerald, Garrick and Neale to Strand Moraines to get erratic samples and visit K2. Excellent travel on the Bowers Piedmont Glacier. Fink and Iles arrived Butter Point from Scott Base, late evening in Snow Trac 35 to join K14.
11 Survey Beacons grid established off Butter Point by the whole party using toboggans.
12 The morning and early afternoon was spent on toboggan repairs after a helo drop of spare parts. One hole was attempted and a partial core recovered, 5km from Butter Point on a line to Hjorth Hill.
13 Surveyors (Fink, Neale) and Garrick to Trig Herb, re-establish the trig cairn. Ward, Pyne, Fitzgerald, Iles, Fink and Neale to Ferrar Glacier Snout, where coring attempt was unsuccessful. Travel was very slow on bare "rough pancake" sea ice. Return via snow-covered Bowers Piedmont for easier travel in early morning.
14 Broke Camp and departed Butter Point for Cape Chocolate (afternoon), arriving early evening near the Garwood Valley where camp was established on the sea ice.
15 Ward, Pyne, Fitzgerald and Iles depart for Scott Base on the Toboggans in marginal visibility conditions. Route to the Daily Islands - "Dirty Ice" - Scott Base was excellent travelling, on bare smooth sea ice until turning towards McMurdo at tip of the "Dirty Ice". The surveyors (Fink, Neale = K3) and Garrick remain for 24 hour tide levelling survey.
16 Rest day at Scott Base. K3 and Garrick return to Base.
17-18 Toboggans and corer repaired.
19 Successfully blasted the ice quarry, for the base emergency water supply, with the help of several Scott Base staff.
20 Remaining repairs and modification to the [unclear: core] trigger carried out.
21 Departed Scott Base for Butter Point (Ward, Pyne, Iles and Garrick). Trip aborted when a toboggan broke down 20km from McMurdo Station, party returned to base overnight for new parts. Fitzgerald remaining at Scott Base for final exam.
22 Pyne and Iles on toboggans return to abandoned toboggan and sledges. Ward and Garrick by helo to sledges, but diverted to Winke Drill site in the lower Taylor Valley after failing to locate the sledges from the air. Return by helo to sledges early afternoon and the party proceeded to Butter Point, arriving early evening.
23 Tent day, bad weather.
24 Coring attempt unsuccessful. Pyne and Iles return to Scott Base page 62 with the corer on an "opportunity" helo.
25 Departed Scott Base early evening in Sno Trac 35, (Pyne, Fitzgerald, Dibble, Iles). Arrived Butter Point early morning (25th) with orange-peel grab borrowed from the Bio. Lab. at McMurdo.
26 Seismic programme began by Dibble and Iles and sampling was continued with orange-peel grab in New Harbour.
27 Completion of the sampling line to Hjorth hill. Seismic programme continued.
28 Repaired sledges and regular maintenance for toboggans. Seismic continued.
29 Tent day. Seismic continued.
30 Completed sampling in the Taylor Valley part of New Harbour and visited the Winkie drill site. Seismic continued.
Dec. 1 Tent day.
2 Completed Ferrar leg of sediment sampling (late evening and early morning of 3rd). Seismic continued. K3 arrived and pitched camp in the late evening.
3 K3 use two toboggans for ice movement survey. Seismic programme came to the end as Sno Trac 35 broke down and party awaiting helo with Bruce Scott (Base mechanic) and Sno Trac parts.
4 Early morning departure for Scott Base, (Ward, Pyne, Dibble, Fink and Neale) with three toboggans and sledges. Excellent ice conditions taking direct route and arrived at 0630. Quick 5.5 hour trip. Scott, Garrick, Fitzgerald and Iles arrive at Base with Sno Trac 35 at 1730 hours.
5 Processed sediment samples, and packed for shipment to N.Z. Unpacked and returned some field equipment.
6 Completed sampling at the McMurdo Desalination Effluent Discharge sites.
7 Rest day.
8-9 Cleaned equipment and packed cargons. Iles by helo to K13 at Knobhead (9th).
10 Prepared equipment for Lashly Mountains. Dibble with K5 to Erebus.
11 Pyne, Ward, Fitzgerald and Garrick to Lashly Mountains, St. Crean, established camp and began work on section L2.
12-14 Continued detailed work on the Weller Coal Measures (L2).
15 Tent day awaiting return to Scott Base, Helo attempt in early evening aborted because of snow storms.
16 Departed Lashly Mountains 1130 hours swapping tents with K13 who finally arrived on the incoming helo.
17 Packed remaining equipment for shipment to N.Z. and returned field equipment to storeman.
18-20 Awaiting C130 flight to Christchurch.
20 Late evening departure for Christchurch (Ward, Pyne, Fitzgerald and Garrick).
page 63